# Evidence of isospin-symmetry violation in high-energy collisions of atomic nuclei – Supplementary Information

### v3

@@ 5 Charge-uniform ensembles and kaons
However, this test is almost impossible to perform experimentally. Neither neutron beams nor neutron targets are easy to produce. Measurements of $K^{0}$ production are possible only through its strong interactions with the detector material. It is also a challenging task. In view of these difficulties and based on previous consideration [Gazdzicki:1991ih], NA61/SHINE performed the test of the charge-symmetry invariance in the kaonic sector, measuring the ratio of charged ($K^{+}+K^{-}$) to neutral ($K^{0}+\bar{K}^{0}=2\cdot K^{0}_{S}$) kaons in nucleus-nucleus ($Z\approx N)$ collisions. In this and the following sections, we elaborate on this test method.
@@ 5.1 Conceptual proof
    Assume the ensemble of initial states is invariant under charge-symmetry transformation. It means the ensemble has an equal population of charge-symmetry partners. It is the charge-uniform ensemble. Here, the important example is the ensemble of two colliding nuclei, each with equal numbers of protons $Z$ and neutrons $N$, $Z=N=A/2$. For instance, this can be an ensemble of $^{16}O$ ++ $^{16}O$ states. As discussed in Sec. 3, point (B), each state of the ensemble has $I_{z}=0$ and thus it is invariant under charge transformation. Obviously, the ensemble of these states is charge-uniform. The second example is the ensemble of an equal number of $\pi^{+}+C$ and $\pi^{-}+C$ systems. As discussed in Sec. 3 point (C), the states have $I_{z}=\pm 1$. Thus, the overall ensemble is charge-uniform, despite each initial state, taken separately, not being charge-symmetry invariant. The experimental study of $^{16}O$ ++ $^{16}O$ as well as $\pi^{+}+C$ and $\pi^{-}+C$ collisions was recently initiated by NA61/SHINE [Grebieskow:2941113].
@@
    From the assumptions (i) and (ii), it trivially follows that the ensemble of final states is charge-symmetry invariant (charge uniform). Namely, a charge-symmetric initial ensemble leads to a charge symmetric final ensemble. This means that the mean multiplicities of charge-symmetry partners have to be equal,
@@ 5.2 Analytical proof
Here, we present an alternative proof of Eqs. 11 and 12 based on the quantum-mechanical formalism. For simplicity, we restrict to the A+A collisions with $N=Z/2$.
@@ 5.3 Comments on isospin-uniform ensembles
The conceptual proof (Sec. 5.1) of the relations (11) and (12) shows similarity to the setup considered by Shmushkevich et al. [Shmushkevich:55, Shmushkevich:56] in the mid 1950s. The setup allowed him to formulate an easy method for relating cross-sections for various processes, following from the invariance of strong interactions under isospin transformations; see also Refs. [MacFarlane:1965wp, 1982AmJPh..50..748W, Pal:2014]. The setup assumes that an initial ‘isospin-uniform’ ensemble of hadronic states (that is, one with an equal mean number of each member of any isospin multiplet) evolving into an isospin-uniform final-state ensemble.
Here we illustrate the Shmushkevich setup considering a simple example of $\rho$ ($I=1$) meson decays. If initially we have equal number of $\rho^{+}$, $\rho^{0}$ and $\rho^{-}$ independent states, the corresponding statistical operator reads:
@@
As such, it is invariant under any isospin transformation $\hat{O}=e^{i\theta_{j}\hat{I}_{j}}$ with $\hat{O}\hat{\rho}\hat{O}^{\dagger}=\hat{\rho}$ for an arbitrary choice of the angles $\theta_{j}.$ With this initial ensemble, the multiplicities of the members of any isospin multiplet need to be equal at any time $t$. This point can be shown by considering the isospin transformation $\hat{C}^{1/2}=e^{iI_{y}\pi/2}$, leading to [^5] Under this transformation, $u\rightarrow(u+d)/2$, $d\rightarrow(-u+d)/2$. Note, a number operator for a certain isospin multiplet member $r$ is given by $\hat{N}_{r}=\sum_{\mathbf{p}}a(\mathbf{p},r)^{\dagger}a(\mathbf{p},r)$, with $\mathbf{p}$ being the momentum. Hence, a transformation with arbitrary $\theta_{3}$ but with $\theta_{1,2}=0$ assures that all expectations of mixed products $\langle a(\mathbf{p},r)^{\dagger}a(\mathbf{p},l)\rangle$ vanish for $r\neq l$. Then, under $\hat{C}^{1/2}$, the operator $N_{\pi^{+}}$ leads to Eq. 16.
@@
In a nucleus-nucleus collision, the initial states are, in general, not isospin-uniform, see Sec. 3. Only if both nuclei have vanishing isospin, the initial state has $I=0$ and is trivially uniform [^6] Within our vector meson analogy, the case with $I=0$ corresponds to $\hat{\rho}=\left|\omega(782)\right\rangle\left\langle\omega(782)\right|$, where $\left|\omega(782)\right\rangle$ is an isosinglet state [ParticleDataGroup:2024cfk]. . For example, the ensemble of $I=1$ and $I_{z}=0$ states is not isospin uniform. Namely, the uniformity would require having, besides the $I_{z}=0$ member, also the ones with $I_{z}=\pm 1$, which, however, are not present in the system since only $I_{z}=0$ is allowed for $N=Z$ nuclei. So, we cannot predict that each member of any isospin multiplet is produced equally (for example, one cannot show that $\langle\pi^{+}\rangle=\langle\pi^{0}\rangle$).
Using the $\rho(770)$-meson example, the case encountered in nucleus-nucleus collisions with $Q/B=1/2$ and $I=1$ is analogous to the charge-invariant statistical operator of the initial ensemble, the $\hat{\rho}=\left|\rho^{0}(770)\right\rangle\left\langle\rho^{0}(770)\right|$, where only the $I_{z}=0$ is retained. This implies that the final ensemble of states is also charge-symmetry invariant, hence $\langle\pi^{+}\rangle=\langle\pi^{-}\rangle,$ $\langle K^{+}\rangle=\langle K^{0}\rangle,$ $\langle K^{-}\rangle=\langle\bar{K}^{0}\rangle,$ $\langle p\rangle=\langle n\rangle$, but it does not imply that $\langle\pi^{+}\rangle=\langle\pi^{0}\rangle$.
In summary, the initial ensemble of nucleus-nucleus collisions with an equal number of protons and neutrons is the charge-uniform ensemble of states. It thus leads to the charge-uniform ensemble of final states. In general, these ensembles do not fulfil the requirement of isospin-uniform ensembles requested for the Shmushkevich method to be valid, not even for $Q/B=1/2$. Nevertheless, charge-symmetry invariance of the initial charge-uniform ensemble is sufficient for having the same mean multiplicities of charged and neutral kaons.
@@ 6 Neutral kaons and $R_{K}$
The neutral kaons are eigenstates of strong interaction, oscillating ($K^{0}\leftrightarrow\bar{K}^{0}$) due to weak interaction. The weak eigenstates are the long-lived ‘K-long’ $K_{L}^{0}$ and short-lived ‘K-short’ $K_{S}^{0}$, which decay into three and two pions with mean lifetimes of 15 m/$c$ and 2.8 cm/$c$, respectively. By neglecting at first the small CP violation (see below), the weak and strong kaonic eigenstates are related by:
@@
This prediction serves as the reference value to test isospin symmetry (specifically charge symmetry) using results from the NA61/SHINE experiment [NA61SHINE:2023azp] and previous experiments; see the next section.
Including CP-symmetry breaking into the mixing modifies the mixing matrix as [Thomson:2013zua]:
@@ 7 Violation of charge-symmetry invariance by known effects
As discussed in Secs. 3 and 4, isospin symmetry is an approximate symmetry of strong interactions. This is because the light quark masses, $m_{u}\neq m_{d}$, are different. In addition, the symmetry is explicitly broken by weak and electromagnetic processes. Breaking by Weak processes is negligible, as shown in the previous section. Electromagnetic effects have been discussed in Ref. [NA61SHINE:2023azp]. While simple estimates suggest they are small, more studies will be required in the future.
As pointed out in Sec. 5, the NA61/SHINE method tests the charge-symmetry invariance, which is a part of isospin symmetry. The experimental quantity used in the test is the ratio of charge-to-neutral kaons, $R_{K}$. Here, we extend the previous discussion in [NA61SHINE:2023azp] about the $R_{K}$ sensitivity on effects violating the charge-symmetry invariance within strong interactions. The quantitative calculations are performed within the statistical Hadron Resonance-Gas (HRG) model [Vovchenko:2019pjl]. The section is closed with a discussion of the charge-symmetry-invariance violation in the charm sector.
@@ 7.1 Violation in the kaon sector
Figure 1: Left: $R_{K}$ as function of $\sqrt{s_{NN}}$ calculated within HRG [Vovchenko:2019pjl] for $Q/B=0.5$ for different cases (from down upwards): assuming exact isospin symmetry: $R_{K}=1$ as expected; including PDG kaon masses; including PDG $\phi(1020)$ branching ratios; including PDG data for all particles. Right: as in the left panel, but for $Q/B=0.4$.
Figure 2: Left: $R_{K}$ as function of $\sqrt{s_{NN}}$ calculated within HRG [Vovchenko:2019pjl] for $Q/B=0.5$ setting different branching ratios of $a_{0}(980)$ and $f_{0}(980)$ resonances (from down upwards): $a_{0}(980)$ and $f_{0}(980)$ with isospin symmetric decays; $a_{0}(980)$ and $f_{0}(980)$ with realistic isospin breaking branching ratios; $a_{0}(980)$ and $f_{0}(980)$ with $K\bar{K}$ mode going entirely into $K^{+}K^{-}$ (unrealistic extreme choice). Right: as in the left panel, but for $Q/B=0.4$.
Different $u$ and $d$ quark masses violate $ud$-flavour symmetry and consequently charge-symmetry. As the quark-mass difference with respect to the QCD mass scale $\Lambda_{QCD}$ is small [Kneur:2011vi, Deur:2016tte], the ratio $R_{K}$ can be expanded as:
$R_{K}=R_{K}^{(0)}+\frac{m_{d}-m_{u}}{\Lambda_{QCD}}R_{K}^{(1)}+$$\text{ ,}$ (27)
where $R_{K}^{(0)}$ is the value of the ratio for the isospin symmetry ($m_{d}=m_{u}$) being equal to one for $Q/B=0.5$. The question is whether $R_{K}^{(1)}+$... terms can reproduce the experimental data. We list below the considered effects and quantify their impact on $R_{K}$ using the HRG model [Vovchenko:2019pjl]. In Ref. [NA61SHINE:2023azp], the HRG results were cross-checked within the microscopic, UrQMD transport model [Bass:1998ca, Bleicher:1999xi, Bleicher:2022kcu].
@@
    In Fig. 1 (left), we have verified that for $Q/B=0.5$ the ratio $R_{K}=1$ is reproduced in the isospin-symmetric limit by the HRG. For this test, the masses of hadrons, resonances, and their decays were set to values that exactly obey isospin symmetry. For $Q/B=0.4$ (the right panel), the isospin-symmetric HRG predicts an increase in $R_{K}$ with collision energy that reaches one above 10 GeV. More neutrons than protons (more $d$ than $u$ quarks) in the initial ensemble favours neutral kaon production. The effect of the initial ensemble asymmetry diminishes at high energies because of the dominance of created up and down quarks over the constituent quarks.
@@
    In connection to the previous point, other potentially relevant $\bar{K}K$ decays refer to the resonances $a_{0}(980)$ and $f_{0}(980)$, whose masses are close to the kaon-kaon threshold. In the PDG table for both resonances, only $\bar{K}K$ is quoted. The standard HRG model has the same branching ratios to $K^{+}K^{-}$ and $K^{0}\bar{K}^{0}$ channels. In particular, in the HRG used in this work, $a_{0}^{0}(980)$ has two decay channels: $\pi^{0}\eta$ ($85\%$) and $K\bar{K}$ ($15\%$), with $7.5\%$ for $K^{+}K^{-}$ and $7.5\%$ for $K^{0}\bar{K}^{0}$. The resonance $f_{0}(980)$ has in HRG the decay channels $\pi^{+}\pi^{-}$ ($52.7\%$), $\pi^{0}\pi^{0}$ ($26.3\%$), and $K\bar{K}$ ($21\%$), with $10.5\%$ for $K^{+}K^{-}$ and $10.5\%$ for $K^{0}\bar{K}^{0}$. Yet charge-symmetry-breaking may be relevant for these resonances. Using Flatté-like spectral functions as in Refs. [Flatte:1976xv, Baru:2004xg, Giacosa:2021mbz] and taking into account the difference of the charged and neutral kaon masses, we find for both resonances a charged-to-neutral ratio of decay widths of about $K^{+}K^{-}/K^{0}\bar{K}^{0}\simeq 1.1$ This ratio remains smaller than $1.2$ when the parameters of the Flatté-like functions are varied within reasonable ranges (see e.g. the compilation of couplings in Ref. [Wu:2008hx]). We thus consider the HRG approach by assigning $20\%$ more charged over neutral kaons for both resonances $a_{0}^{0}(980)$ and $f_{0}(980)$. In all cases, the ratio $R_{K}$ changes by $0.36\%$ at large $\sqrt{s_{NN}}$ . We also considered the ‘extreme scenario’ where $a_{0}^{0}(980)$ and $f_{0}(980)$ decay only into charged kaons. It gives an upper limit of the impact of $a_{0}^{0}(980)$ and $f_{0}(980)$ decays on $R_{K}$. There is no significant effect for $\sqrt{s_{NN}}\lesssim 10$ GeV, and a slight (1.84%) increase of $R_{K}$ at high energies, see Fig. 2.

### v4

@@ 5 Charge-uniform ensembles and kaons
However, this test is almost impossible to perform experimentally. Neither neutron beams nor neutron targets are easy to produce. Measurements of $K^{0}$ production are possible only through its strong interactions with the detector material. It is also a challenging task. In view of these difficulties and based on previous consideration [22], NA61/SHINE performed the test of the charge-symmetry invariance in the kaonic sector, measuring the ratio of charged ($K^{+}+K^{-}$) to neutral ($K^{0}+\bar{K}^{0}=2\cdot K^{0}_{S}$) kaons in nucleus-nucleus ($Z\approx N)$ collisions. In this and the following sections, we elaborate on this test method.
@@ 5.1 Conceptual proof
    Assume the ensemble of initial states is invariant under charge-symmetry transformation. It means the ensemble has an equal population of charge-symmetry partners. It is the charge-uniform ensemble. Here, the important example is the ensemble of two colliding nuclei, each with equal numbers of protons $Z$ and neutrons $N$, $Z=N=A/2$. For instance, this can be an ensemble of $^{16}O$ ++ $^{16}O$ states. As discussed in Sec. 3, point (B), each state of the ensemble has $I_{z}=0$, and thus it is invariant under charge transformation. Obviously, the ensemble of these states is charge-uniform. The second example is the ensemble of an equal number of $\pi^{+}+C$ and $\pi^{-}+C$ systems. As discussed in Sec. 3 point (C), the states have $I_{z}=\pm 1$. Thus, the overall ensemble is charge-uniform, despite each initial state, taken separately, not being charge-symmetry invariant. The experimental study of $^{16}O$ ++ $^{16}O$ as well as $\pi^{+}+C$ and $\pi^{-}+C$ collisions was recently initiated by NA61/SHINE [31].
@@
    From the assumptions (i) and (ii), it trivially follows that the ensemble of final states is charge-symmetry invariant (charge uniform). A charge-symmetric initial ensemble yields a charge-symmetric final ensemble. This means that the mean multiplicities of charge-symmetry partners have to be equal,
@@ 5.2 Analytical proof
Here, we present an alternative proof of Eqs. 11 and 12 based on the quantum-mechanical formalism. For simplicity, we restrict to the A+A collisions with $N=Z=A/2$.
@@ 5.3 Comments on isospin-uniform ensembles
The conceptual proof (Sec. 5.1) of the relations (11) and (12) shows similarity to the setup considered by Shmushkevich [51, 17] in the mid 1950s. The setup allowed him to formulate an easy method for relating cross-sections for various processes, following from the invariance of strong interactions under isospin transformations; see also Refs. [43, 58, 47]. The main idea is that an initial ‘isospin-uniform’ ensemble of hadronic states (that is, one with an equal mean number of each member of any isospin multiplet) evolves into an isospin-uniform final-state ensemble.
Here we illustrate the Shmushkevich method considering a simple example of $\rho(770)$ ($I=1$) meson decays. If initially we have equal number of $\rho(770)^{+}$, $\rho(770)^{0}$ and $\rho(770)^{-}$ independent states, the corresponding statistical operator reads:
@@
As such, it is invariant under any isospin transformation $\hat{O}=e^{i\theta_{j}\hat{I}_{j}}$ with $\hat{O}\hat{\rho}\hat{O}^{\dagger}=\hat{\rho}$ for an arbitrary choice of the angles $\theta_{j}.$ With this initial ensemble, the multiplicities of the members of any isospin multiplet must be equal at all times $t$. This point can be shown by considering the isospin transformation $\hat{C}^{1/2}=e^{iI_{y}\pi/2}$, leading to [^5] Under this transformation, $u\rightarrow(u+d)/\sqrt{2}$, $d\rightarrow(-u+d)/\sqrt{2}$. Note, a number operator for a certain isospin multiplet member $r$ is given by $\hat{N}_{r}=\sum_{\mathbf{p}}a(\mathbf{p},r)^{\dagger}a(\mathbf{p},r)$, with $\mathbf{p}$ being the momentum. Hence, a transformation with arbitrary $\theta_{3}$ but with $\theta_{1,2}=0$ assures that all expectations of mixed products $\langle a(\mathbf{p},r)^{\dagger}a(\mathbf{p},l)\rangle$ vanish for $r\neq l$. Then, under $\hat{C}^{1/2}$, the operator $N_{\pi^{+}}$ leads to Eq. 16.
@@
In a nucleus-nucleus collision, the initial states are, in general, not isospin-uniform, see Sec. 3. Only if both nuclei have vanishing isospin, the initial state has $I=0$ and is trivially uniform [^6] Within our vector meson analogy, the case with $I=0$ corresponds to $\hat{\rho}=\left|\omega(782)\right\rangle\left\langle\omega(782)\right|$, where $\left|\omega(782)\right\rangle$ is an isosinglet state [46]. . For example, the ensemble of $I=1$ and $I_{z}=0$ states is not isospin-uniform. Namely, the uniformity would require having, besides the $I_{z}=0$ member, also the ones with $I_{z}=\pm 1$, which, however, are not present in the system since only $I_{z}=0$ is allowed for $N=Z$ nuclei. Thus, we cannot predict that each member of an isospin multiplet is produced with equal probability (for example, one cannot show that $\langle\pi^{+}\rangle=\langle\pi^{0}\rangle$).
Using the $\rho(770)$-meson example, the case encountered in nucleus-nucleus collisions with $Q/B=1/2$ and $I=1$ is analogous to the charge-invariant statistical operator of the initial ensemble, the $\hat{\rho}=\left|\rho^{0}(770)\right\rangle\left\langle\rho^{0}(770)\right|$, where only the member with $I_{z}=0$ is retained. This implies that the final ensemble of states is also charge-symmetry invariant, hence $\langle\pi^{+}\rangle=\langle\pi^{-}\rangle,$ $\langle K^{+}\rangle=\langle K^{0}\rangle,$ $\langle K^{-}\rangle=\langle\bar{K}^{0}\rangle,$ $\langle p\rangle=\langle n\rangle$, but it does not imply that $\langle\pi^{+}\rangle=\langle\pi^{0}\rangle$. (In fact, $\rho^{0}(770)$ decays solely into $\pi^{+}\pi^{-}$, implying no $\pi^{0}$ production.)
In summary, the initial ensemble of nucleus-nucleus collisions with an equal number of protons and neutrons is the charge-uniform ensemble of states. It thus leads to the charge-uniform ensemble of final states. In general, these ensembles do not fulfil the requirement of isospin-uniform ensembles that is necessary for the Shmushkevich method to be valid, not even for $Q/B=1/2$. Nevertheless, charge-symmetry invariance of the initial charge-uniform ensemble is sufficient for having the same mean multiplicities of charged and neutral kaons.
@@ 6 Neutral kaons and $R_{K}$
The neutral kaons are eigenstates of the strong interaction and oscillate ($K^{0}\leftrightarrow\bar{K}^{0}$) due to the weak interaction. The weak eigenstates are the long-lived ‘K-long’ $K_{L}^{0}$ and short-lived ‘K-short’ $K_{S}^{0}$, which decay into three and two pions with mean lifetimes of $15\penalty 10000\ m/c$ and 2.8 cm/$c$, respectively. By neglecting at first the small CP violation (see below), the weak and strong kaonic eigenstates are related by:
@@
This prediction serves as the reference value for testing isospin symmetry (specifically, charge symmetry) using results from the NA61/SHINE experiment [6] and previous experiments; see the next section.
Including CP-symmetry breaking into the mixing modifies the mixing matrix as [52]:
@@
This modification due to weak interaction can be safely neglected.
@@ 7 Violation of charge-symmetry invariance by known effects
As discussed in Secs. 3 and 4, isospin symmetry is an approximate symmetry of strong interactions. This is because the light quark masses, $m_{u}\neq m_{d}$, are different. In addition, the symmetry is explicitly broken by weak and electromagnetic processes. Breaking by weak processes is negligible, as shown in the previous section. Electromagnetic effects have been discussed in Ref. [6]. While simple estimates suggest they are small, further studies will be required.
As pointed out in Sec. 5, the NA61/SHINE method tests the charge-symmetry invariance, which is a part of isospin symmetry. The experimental quantity used in the test is the ratio of charge-to-neutral kaons, $R_{K}$. Here, we extend the previous discussion in [6] about the $R_{K}$ sensitivity on effects violating the charge-symmetry invariance within strong interactions. The quantitative calculations are performed within the statistical Hadron Resonance Gas model [54]. In doing so, also the fact that $Q/B<1/2$ must be taken into account. For Ar-Sc used in NA61/SHINE, $Q/B=0.459$, while for lead-lead scattering $Q/B=0.394$. Thus, the choice $Q/B=0.4$ shall also be investigated in HRG calculations.
The section concludes with a discussion of charge-symmetry-invariance violation in the charm sector.
@@ 7.1 Violation in the kaon sector
Figure 2: Left: $R_{K}$ as a function of $\sqrt{s_{NN}}$ calculated within HRG [54] for $Q/B=0.5$ for different cases (from down upwards): assuming exact isospin symmetry: $R_{K}=1$ as expected; including PDG kaon masses; including PDG $\phi(1020)$ branching ratios; including PDG data for all particles. Right: as in the left panel, but for $Q/B=0.4$.
Figure 3: Left: $R_{K}$ as a function of $\sqrt{s_{NN}}$ calculated within HRG [54] for $Q/B=0.5$ setting different branching ratios of $a_{0}(980)$ and $f_{0}(980)$ resonances (from down upwards): $a_{0}(980)$ and $f_{0}(980)$ with isospin symmetric decays; $a_{0}(980)$ and $f_{0}(980)$ with realistic isospin breaking branching ratios; $a_{0}(980)$ and $f_{0}(980)$ with $K\bar{K}$ mode going entirely into $K^{+}K^{-}$ (unrealistic extreme choice). Right: as in the left panel, but for $Q/B=0.4$.
Different $u$ and $d$ quark masses violate $ud$-flavour symmetry and consequently charge-symmetry. We list below the considered effects and quantify their impact on $R_{K}$ using the HRG model [54]. In Ref. [6], the HRG results were cross-checked within the microscopic, UrQMD transport model [10, 12, 13].
@@
    In Fig. 2 (left), we have verified that for $Q/B=0.5$ the ratio $R_{K}=1$ is reproduced in the isospin-symmetric limit by the HRG. For this test, the masses of hadrons, resonances, and their decays were set to values that exactly obey isospin symmetry. For $Q/B=0.4$ (the right panel), the isospin-symmetric HRG predicts an increase in $R_{K}$ with collision energy that reaches one above 10 GeV. More neutrons than protons (more $d$ than $u$ quarks) in the initial ensemble favours neutral kaon production. The effect of the initial ensemble asymmetry diminishes at high energies because of the dominance of created up and down quarks over the constituent quarks. In this limit, the theoretical predictions are very similar to the $Q/B=1/2$ case.
@@
    In connection to the previous point, other potentially relevant $\bar{K}K$ decays refer to the resonances $a_{0}(980)$ and $f_{0}(980)$, whose masses are close to the kaon-kaon threshold. In the PDG table for both resonances, only $\bar{K}K$ is quoted. The standard HRG model has the same branching ratios to $K^{+}K^{-}$ and $K^{0}\bar{K}^{0}$ channels. In particular, in the HRG used in this work, $a_{0}^{0}(980)$ has two decay channels: $\pi^{0}\eta$ ($85\%$) and $K\bar{K}$ ($15\%$), with $7.5\%$ for $K^{+}K^{-}$ and $7.5\%$ for $K^{0}\bar{K}^{0}$. The resonance $f_{0}(980)$ has in HRG the decay channels $\pi^{+}\pi^{-}$ ($52.7\%$), $\pi^{0}\pi^{0}$ ($26.3\%$), and $K\bar{K}$ ($21\%$), with $10.5\%$ for $K^{+}K^{-}$ and $10.5\%$ for $K^{0}\bar{K}^{0}$. Yet charge symmetry violation may be relevant for these resonances. Using Flatté-like spectral functions as in Refs. [19, 9, 28] and taking into account the difference of the charged and neutral kaon masses, we find for both resonances a charged-to-neutral ratio of decay widths of about $K^{+}K^{-}/K^{0}\bar{K}^{0}\simeq 1.1$. This ratio remains smaller than $1.2$ when the parameters of the Flatté-like functions are varied within reasonable ranges (see e.g. the compilation of couplings in Ref. [59]). We thus consider the HRG approach by assigning $20\%$ more charged over neutral kaons for both resonances $a_{0}^{0}(980)$ and $f_{0}(980)$. In all cases, the ratio $R_{K}$ changes by $0.36\%$ at large $\sqrt{s_{NN}}$ . We also considered the ‘extreme scenario’ where $a_{0}^{0}(980)$ and $f_{0}(980)$ decay only into charged kaons. It gives an upper limit of the impact of $a_{0}^{0}(980)$ and $f_{0}(980)$ decays on $R_{K}$. There is no significant effect for $\sqrt{s_{NN}}\lesssim 10$ GeV, and a slight (1.84%) increase of $R_{K}$ at high energies, see Fig. 3.